OpenStudy (anonymous):

Solve the matrix equation. [2,1,4,3][x,y]=[10,-2]

OpenStudy (e.mccormick):

This?\[\left[\begin{matrix} 2 & 1\\ 4 & 3 \end{matrix}\right] \left[\begin{matrix} x\\ y \end{matrix}\right]=\left[\begin{matrix} 10\\ -2 \end{matrix}\right]\]

OpenStudy (e.mccormick):

There are a few ways to solve a matrix. Are they saying you need to use a specific method, or just to get the answer?

OpenStudy (e.mccormick):

Cramer's Rule \[x=\frac{|A_x|}{|A|}\quad x=\frac{|A_y|}{|A|}\] Gauss Row reduction until it is in upper triangular form. Then use back substitution. Gauss-Jordan \[Ax=b\implies Is=x\]where s is a solution found through reduced row echelon form. If you still need help when you get back online, reply and I (or many others) can work with you on any of those.

OpenStudy (e.mccormick):

oops... and that second x in Cramer's is supposed to be a y.

OpenStudy (anonymous):

Yes that's the equation I meant

OpenStudy (anonymous):

there are many methods, so you'd need to let me know which to use. you are probably using either row reduction or using the inverse. lmk :)

OpenStudy (e.mccormick):

\[\left[\begin{matrix} 2 & 1\\ 4 & 3 \end{matrix}\right] \left[\begin{matrix} x\\ y \end{matrix}\right]=\left[\begin{matrix} 10\\ -2 \end{matrix}\right]\implies \left[\begin{matrix} 2 & 1&|&10\\ 4 & 3&|&-2 \end{matrix}\right] \]That is the setup for row reduction. If you are using that, that is the start. If you are being told to use Cramers: \[A=\left[\begin{matrix} 2 & 1\\ 4 & 3 \end{matrix}\right],\quad A_x=\left[\begin{matrix} 10 & 1\\ -2 & 3 \end{matrix}\right],\quad A_y=\left[\begin{matrix} 2 & 10\\ 4 & -2 \end{matrix}\right] \]Then solve: \[x=\frac{|A_x|}{|A|}\quad y=\frac{|A_y|}{|A|}\]Where the |A| stuff means determinant.